Solve the Lagrange’s Linear Equation (y^2+z^2-x^2)p-2xyq+2xz=0
1
Expert's answer
2020-08-09T15:21:32-0400

"(y^2+z^2-x^2)p-2xyq+2xz=0\\
\frac{dx}{y^2+z^2-x^2}=\frac{dy}{-2xy}=\frac{dz}{-2xz}\\
\frac{dy}{-2xy}=\frac{dz}{-2xz}\\
\frac{dy}{y}=\frac{dz}{z}\\
\log y=\log z+\log C_1\\
C_1=\frac{y}{z}\\
\text{each fraction}=\frac{xdx+ydy+zdz}{x(y^2+z^2-x^2)-2xy^2-2xz^2}=\\
=\frac{xdx+ydy+zdz}{-x(y^2+z^2+x^2)}\\
\frac{xdx+ydy+zdz}{-x(y^2+z^2+x^2)}=\frac{dy}{-2xy}\\
\frac{2xdx+2ydy+2zdz}{y^2+z^2+x^2}=\frac{dy}{y}\\
\log (y^2+z^2+x^2)=\log y+\log C_2\\
C_2=\frac{y^2+z^2+x^2}{y}\\
f(C_1, C_2)=0\\
f(\frac{y}{z}, \frac{y^2+z^2+x^2}{y})=0"